# Beyond Text: Utilizing Vocal Cues to Improve Decision Making in LLMs for Robot Navigation Tasks

### v2

@@ Abstract
While LLMs excel in processing text in these human conversations, they struggle with the nuances of verbal instructions in scenarios like social navigation, where ambiguity and uncertainty can erode trust in robotic and other AI systems. We can address this shortcoming by moving beyond text and additionally focusing on the paralinguistic features of these audio responses. These features are the aspects of spoken communication that do not involve the literal wording (lexical content) but convey meaning and nuance through how something is said. We present Beyond Text; an approach that improves LLM decision-making by integrating audio transcription along with a subsection of these features, which focus on the affect and more relevant in human-robot conversations.This approach not only achieves a 70.26% winning rate, outperforming existing LLMs by 22.16% to 48.30% (gemini-1.5-pro and gpt-3.5 respectively), but also enhances robustness against token manipulation adversarial attacks, highlighted by a 22.44% less decrease ratio than the text-only language model in winning rate. “Beyond Text” marks an advancement in social robot navigation and broader Human-Robot interactions, seamlessly integrating text-based guidance with human-audio-informed language models.
@@ 1 Introduction
    A New Dataset on Disfluent Human Audio-Guided Instructions. In support of our framework and to catalyze future research, we collect a dataset comprising 500 human audio navigational instructions with perceptual “ground truth” captured human labeling. This dataset is characterized by a diverse range of semantic and vocal ambiguities, offering an extensive experimental foundation for our proposed method and subsequent investigative endeavors in this space.
@@ 2 Related Works
Large Language Models for Robotics. Recent studies have demonstrated the versatility of LLMs in various robotic contexts. These include employing few-shot summarizing skills for task planning in robots (Wu et al., 2023; Liu et al., 2023b), interpreting contextual information for robotic locomotion decision-making (Shek et al., 2023), integrating visual feedback for open-world robotic manipulation (Lin et al., 2023a; Ahn et al., 2022), and enhancing motion planning in autonomous driving (Mao et al., 2023). For human-robot interaction, Shah’s research Shah et al. (2023) suggests using LLMs for semantic prediction of object locations, thereby optimizing navigation. SayCan Ahn et al. (2022) introduces multimodal input to generate the optimal long-horizon instructions and effectively score the affordance probability of the instruction for robotic task planning. Meanwhile, Ren et al. (2023) develops a framework to assess and synchronize the uncertainty levels in LLM-based planning systems, enabling them to seek assistance when necessary. However, unlike our framework, KnowNo Ren et al. (2023) does not account for potentially unreliable human input but rather relies on human selection among generated options. Beyond Text uniquely concentrates on assessing human uncertainty through affective computing for robotic navigation, particularly when guided by human audio that might be ambiguous or incorrect. This approach allows robots to gauge the reliability of human-audio navigation instructions by analyzing vocal emotions, enabling more informed reasoning.
@@
Foundation Models with Audio. The ability to perceive and interpret auditory information is essential for AI agents in the real world. GPT-4’s recent updateAchiam et al. (2023), incorporating Whisper’s audio-to-text feature (Radford et al., 2023), significantly advances foundational models’ capacity to transcribe audio inputs. However, this transcription process often overlooks nuanced vocal features inherent in speech. Parallel research efforts (Ghosal et al., 2023; Borsos et al., 2023) explore language models for generating audio from text and enhancing audio-visual understanding in videos. Notably, SALMONN Tang et al. (2023) integrates speech audio encoders with a multimodal language model, enabling direct processing of audio inputs for tasks like speech translation, answering spoken queries, etc. Despite these advancements, current research does not specifically address the use of audio inputs in aiding human-robot navigation or in extracting vocal uncertainty from speech.
@@
Uncertainty Modeling in LLMs. Gauging the trustworthiness of responses generated by LLMs remains an open challenge in today’s LLM era, with limited research on uncertainty quantification for natural language generation. Lin Lin et al. (2023b) compared multiple uncertainty measurements in the black-box language model, and Kuhn Kuhn et al. (2023) introduced entropy as a method to model uncertainty in the large language model. In our approach, we take inspiration from prior works (Kuhn et al., 2023; Lin et al., 2023b) and define a confidence score $\mathcal{C}(\rho)$ that is inversely proportional to the Kullback-Leibler divergence between the current LLM’s prediction and ground truth human perception distribution to show the LLM’s confidence on uncertainty measurement.
@@
Human Affective Analysis in Robotics. For social robots to effectively coexist and interact with humans, it is imperative that they comprehend human emotional states for decision-making processes. Emotion understanding from speech for human-robot interaction has been studied in (Lakomkin et al., 2018a), while deep reinforcement learning methods (Dorbala et al., 2021) and cognition model (Eppe et al., 2016) has been used to understand textual ambiguities in natural languages. RxR dataset (Ku et al., 2020) is a well-known audio dataset for visual language navigation but does not cover the affective cues embedded in the vocal instructions, which hold substantial information that can reveal human uncertainty and trust. Beyond Text introduces an audio-augmentation technique for large language models (LLMs) that analyzes textual and vocal cues in human audio commands. This approach enables more human-like robotic systems for social navigation by identifying emotions and ambiguities through both speech content and vocal tones.
@@ 4 New Dataset: Disfluent Navigational Instruction Audio Dataset (DNIA)
To mimic real-world Human-Robot Interaction (HRI) scenarios as closely as possible, we align with the methodologies suggested in LLM evaluation literature (Bang et al., 2023; Chang et al., 2023). We employ human annotation as one of the primary assessment techniques. This involves manual evaluation, which is particularly relevant in HRI contexts due to its ability to capture the nuanced complexities of human interaction. We use human evaluation to calculate the winning rate of analyzing human navigational guidance uncertainty in Section 5.4. Our approach entails labeling human audio-guided instructions through a selection process from a set of five multiple-choice options designed to represent the uncertainty inherent in the content accurately. We recruit a different set of 20 participants with diverse genders and dialect backgrounds to ensure a broad range of perspectives. We developed a user study website (see Appendix G) where these participants listened to navigational audio clips. The instructions emphasized careful listening and unbiased selection, with a constraint of a maximum of one minute per clip to simulate real-world decision-making time pressures.
@@ 5 Experimental Results
To validate the efficacy of our approach, we utilize confidence score and winning rate to show that Beyond Text can achieve low variance and low bias results compared to current LLM-based human instructions without vocal cues. We will first introduce the evaluation metric, and the comparison experiments, ablation studies on various vocal cues and LLM types.
@@ 5.1 Evaluation Metric: Confidence Measure
In this paper, we specifically focus on improving the uncertainty or the confidence score of the LM $p_{\theta}$ on the optimal (ground-truth) candidate responses $y^{j*}$. Specifically, we represent the probability distribution over the possible candidates $\{y^{j}\}_{j=1}^{J}$ by the current LM $p_{\theta}$ as $\rho=[\rho(y^{1}),\rho(y^{2}),\cdots,\rho(y^{J})]$ and the ground-truth be represented as a dirac-delta over the true response as $\rho^{*}=[0,1\cdots 0]$, assuming the second candidate to be the optimal $y^{j*}$ with $j*=2$ without loss of generality. We define a confidence measure $\mathcal{C}(\rho)$ inversely proportional to the Kullback–Leibler divergence between $\rho$ and $\rho^{*}$ given by
$\mathcal{C}(\rho)=\frac{1}{\text{KL}(\rho,\rho^{*})},$ (3)
where we note that the confidence measure will reduce as the KL divergence between $\rho$ and $\rho^{*}$ increases. The range of a confidence measure, which is inversely proportional to entropy, is bounded between $[0,\infty]$. Ideally, we would like to have the confidence measure as high as possible, which requires us to reduce $\text{KL}(\rho,\rho^{*})$. We first expand and note that
$\text{KL}(\rho,\rho^{*})=-H(\rho)=-\sum_{j=1}^{J}\rho(y^{j})\log(\rho^{*}{(y^{j})}),$ (4)
where $H(\rho)=-\sum_{j=1}^{J}\rho(y^{j})\log(\rho{(y^{j})})$ is the entropy of predicted distribution $\rho(y^{j})$ by the LLM which quantifies the amount of information contained in random variables. We note that since we don’t have any control over the second term in equation 4 because $\rho^{*}$ is unknown, we can focus on minimizing the first term in equation 4, which tells us to maximize the entropy in order to reduce the KL divergence between $\rho$ and $\rho^{*}$. We hypothesize that one of the major reasons for this uncertainty is the information loss while considering only textual information $x$ and losing various critical attributes of the input audio information $v$. From the standard information theory, one can quantify this loss in information as $\mathcal{I}_{l}=H(V)-I(X;V)$ as the difference between the entropy of the original audio input random variable $V$ and the mutual information between the text random variable $X$ and $V$ represented as $I(Y;V)$ which measures the amount of information that is shared between $X,V$. Notably, a higher confidence score—which indicates a lower KL divergence between $\rho$ and $\rho^{*}$—is desirable, as it signifies a closer alignment with the ground truth. In this work, we empirically show that adding the vocal cue in the input helps to reduce the $\text{KL}(\rho,\rho^{*})$ by plotting the confidence measure equation 3 for robotic navigation tasks in Section 5.
@@ 5.4 Winning Rate
Besides indicating low variances and multiple choice distribution change, we also calculate the winning rate. We compare the human annotations for each DNIA data clip with the highest log-probability choices generated by the model to ascertain whether the choices and reasoning processes generated by the LLMs are in harmony with human conceptual understanding. The results, as detailed in Table 3, show the limitation of current LLMs in explaining vocal uncertainty from only textual input and the effectiveness of our audio augmentation approach. Given that the audio clips originated from human speakers, the most reliable measure of their trustworthiness was human perception, hence the incorporation of a user study to gauge the winning rate with which the language model identified uncertainties in the audio clips and selected the appropriate next-step action.
@@ Appendix E DNIA Dataset Details and Samples
For VU clips, we ensure no semantic uncertainty signals, such as uncertain-related words (i.e., ‘maybe,’ ‘I guess’), speech repairs, and hesitations. We believe those are common affective characteristics associated with human speech. However, those affective cues are not prominent in LU clips. As our position states that affective cues (duration, pitch, loudness) are associated with speech uncertainty levels in robot navigation tasks, we believe the setup of our dataset is comprehensive enough to show that vocal cues are good predictors. For review purposes, we create an anonymous GitHub repo: https://anonymous.4open.science/r/DNIA-dataset-D22E. The full dataset will be publicly released soon after the review as we prepare them for storage and access control. The dataset will come with textual annotations (grounded labels) as well as Whisper-generated transcription to save resources for other researchers who can reproduce the result.
@@ Appendix F Pitch and Loudness Shifts Detection
0:  audio $\mathcal{V}$, $\theta_{l}$ (loudness threshold), $\theta_{p}$ (pitch threshold)
0:  Timestamps of maximum loudness change ($t_{l}$) and pitch shift ($t_{p}$)
  $L\leftarrow\emptyset$, $P\leftarrow\emptyset$
  $l_{prev},p_{prev}\leftarrow$ loudness and pitch at $[t_{0},t_{1}]$
  for each second $t_{i}$ in $\mathcal{V}$ do
     Define segment as $[t_{i},t_{i+1}]$
     $l_{t_{i+1}}\leftarrow$ avg_loudness(segment)
     $p_{t_{i+1}}\leftarrow$ calculate_pitch(segment)
     if $|l_{t_{i+1}}-l_{prev}|>\theta_{l}$ then
        Append ($t_{i+1}$, $|l_{t_{i+1}}-l_{prev}|$) to $L$
     end if
     if $|p_{t_{i+1}}-p_{prev}|>\theta_{p}$ then
        Append ($t_{i+1}$, $|p_{t_{i+1}}-p_{prev}|$) to $P$
     end if
     $l_{prev}\leftarrow l_{t_{i+1}}$, $p_{prev}\leftarrow p_{t_{i+1}}$
  end for
  $t_{l}\leftarrow\text{argmax}_{t}\left\{\delta_{l_{t}}\ \middle|\ (t,\delta_{l_{t}})\in L\right\}$
  $t_{p}\leftarrow\text{argmax}_{t}\left\{\delta_{p_{t}}\ \middle|\ (t,\delta_{p_{t}})\in P\right\}$
Appendix G User Interface for Human Evaluation on DNIA Dataset

### v3

@@ Abstract
While LLMs excel in processing text in these human conversations, they struggle with the nuances of verbal instructions in scenarios like social navigation, where ambiguity and uncertainty can erode trust in robotic and other AI systems. We can address this shortcoming by moving beyond text and additionally focusing on the paralinguistic features of these audio responses. These features are the aspects of spoken communication that do not involve the literal wording (lexical content) but convey meaning and nuance through how something is said. We present Beyond Text; an approach that improves LLM decision-making by integrating audio transcription along with a subsection of these features, which focus on the affect and more relevant in human-robot conversations. This approach not only achieves a 70.26% winning rate, outperforming existing LLMs by 22.16% to 48.30% (gemini-1.5-pro and gpt-3.5 respectively), but also enhances robustness against token manipulation adversarial attacks, highlighted by a 22.44% less decrease ratio than the text-only language model in winning rate. We also present the first dataset on disfluent human audio-guided instructions (https://github.com/beyond-text/DNIA-dataset) for future research in this field. “Beyond Text” marks an advancement in social robot navigation and broader human-robot interactions, seamlessly integrating text-based guidance with human-audio-informed language models.
@@ 1 Introduction
    A New Dataset on Disfluent Human Audio-Guided Instructions. In support of our framework and to catalyze future research, we collect a dataset comprising 500 human audio navigational instructions with perceptual “ground truth” captured human labeling. This dataset (https://github.com/beyond-text/DNIA-dataset) is characterized by a diverse range of semantic and vocal ambiguities, offering an extensive experimental foundation for our proposed method and subsequent investigative endeavors in this space.
@@ 2 Related Works
Large Language Models for Robotics. Recent studies have demonstrated the versatility of LLMs in various robotic contexts. These include employing few-shot summarizing skills for task planning in robots (Wu et al., 2023; Liu et al., 2023b), interpreting contextual information for robotic locomotion decision-making (Shek et al., 2023), integrating visual feedback for open-world robotic manipulation (Lin et al., 2023a; Ahn et al., 2022), and enhancing motion planning in autonomous driving Mao et al. (2023). For human-robot interaction, Shah et al. (2023) suggests using LLMs for semantic prediction of object locations, thereby optimizing navigation. Ahn et al. (2022) introduces multimodal input to generate the optimal long-horizon instructions and effectively score the affordance probability of the instruction for robotic task planning. Meanwhile, Ren et al. (2023) develops a framework to assess and synchronize the uncertainty levels in LLM-based planning systems, enabling them to seek assistance when necessary. However, unlike our framework, Ren et al. (2023) does not account for potentially unreliable human input but rather relies on human selection among generated options. Beyond Text uniquely concentrates on assessing human uncertainty through affective computing for robotic navigation, particularly when guided by human audio that might be ambiguous or incorrect. This approach allows robots to gauge the reliability of human-audio navigation instructions by analyzing vocal emotions, enabling more informed reasoning.
@@
Foundation Models with Audio. The ability to perceive and interpret auditory information is essential for AI agents in the real world. GPT-4’s recent update(Achiam et al., 2023), incorporating Whisper’s audio-to-text feature (Radford et al., 2023), significantly advances foundational models’ capacity to transcribe audio inputs. However, this transcription process often overlooks nuanced vocal features inherent in speech. Ghosal et al. (2023); Borsos et al. (2023) explore language models for generating audio from text and enhancing audio-visual understanding in videos. Notably, Tang et al. (2023) integrates speech audio encoders with a multimodal language model, enabling direct processing of audio inputs for tasks like speech translation, answering spoken queries, etc. Despite these advancements, current research does not specifically address the use of audio inputs in aiding human-robot navigation or in extracting vocal uncertainty from speech.
@@
Uncertainty Modeling in LLMs. Gauging the trustworthiness of responses generated by LLMs remains an open challenge in today’s LLM era, with limited research on uncertainty quantification for natural language generation. Lin et al. (2023b) compared multiple uncertainty measurements in the black-box language model, and Kuhn et al. (2023) introduced entropy as a method to model uncertainty in the large language model. In our approach, we take inspiration from prior works (Kuhn et al., 2023; Lin et al., 2023b) and define a confidence score $\mathcal{C}(\rho)$ that is inversely proportional to the Kullback-Leibler divergence between the current LLM’s prediction and ground truth human perception distribution to show the LLM’s confidence on uncertainty measurement.
@@
Human Affective Analysis in Robotics. For social robots to effectively coexist and interact with humans, it is imperative that they comprehend human emotional states for decision-making processes. Emotion understanding from speech for human-robot interaction has been studied in Lakomkin et al. (2018a), while deep reinforcement learning methods (Dorbala et al., 2021) and cognition model (Eppe et al., 2016) has been used to understand textual ambiguities in natural languages. RxR dataset (Ku et al., 2020) is a well-known audio dataset for visual language navigation but does not cover the affective cues embedded in the vocal instructions, which hold substantial information that can reveal human uncertainty and trust. Beyond Text introduces an audio-augmentation technique for large language models (LLMs) that analyzes textual and vocal cues in human audio commands. This approach enables more human-like robotic systems for social navigation by identifying emotions and ambiguities through both speech content and vocal tones.
@@ 4 New Dataset: Disfluent Navigational Instruction Audio Dataset (DNIA)
To mimic real-world Human-Robot Interaction (HRI) scenarios as closely as possible, we align with the methodologies suggested in LLM evaluation literature (Bang et al., 2023; Chang et al., 2023). We employ human annotation as one of the primary assessment techniques. This involves manual evaluation, which is particularly relevant in HRI contexts due to its ability to capture the nuanced complexities of human interaction. We use human evaluation to calculate the winning rate of analyzing human navigational guidance uncertainty in Section 5.4. Our approach entails labeling human audio-guided instructions through a selection process from a set of five multiple-choice options designed to represent the uncertainty inherent in the content accurately. We recruit a different set of 20 participants with diverse genders and dialect backgrounds to ensure a broad range of perspectives. We developed a user study website (see Appendix H) where these participants listened to navigational audio clips. The instructions emphasized careful listening and unbiased selection, with a constraint of a maximum of one minute per clip to simulate real-world decision-making time pressures.
@@ 5 Experimental Results
To validate the efficacy of our approach, we utilize confidence score and winning rate to show that Beyond Text can achieve low variance and low bias results compared to current LLM-based human instructions without vocal cues. We will first introduce the evaluation metric and the comparison experiments and ablation studies on various vocal cues and LLM types.
@@ 5.1 Evaluation Metric: Confidence measure
In this paper, we specifically focus on improving the uncertainty or the confidence score of the LM $p_{\theta}$ on the optimal (ground-truth) candidate responses $y^{j*}$.
We note that as defined in equation 1, our model is given by $p_{\theta}(\cdot|Q(v))$ where $Q(v)$ is combination of vocal and text cue. For a given $Q(v)$, let us consider $\{y^{j}\}_{j=1}^{J}$ as the set of possible candidates, and the probability of each candidate is given by $p_{\theta}(y^{j}|Q(v))$, and we define the probability distribution $\rho$ as $\rho:=[p_{\theta}(y^{1}|Q(v)),p_{\theta}(y^{2}|Q(v)),\cdots,p_{\theta}(y^{J}|Q(v))]$. Hence, this is clearly interpratable in the sense that it is the probability distribution across possible candidates.
We do not assume that the gound truth distribution $\rho^{*}$ is known or have any specific form but it holds that $\rho^{*}(y^{j})>0$ for all $j$ so that it is a valid distribution and $\sum_{j}\rho^{*}(y^{j})=1$. In most of the practical cases, the ground truth distribution will be concentrated around the true response $y^{j*}$ if it is unique. Now, we define the confidence measure $\mathcal{C}(\rho)$ of a distribution $\rho$ as
$\mathcal{C}(\rho):=\frac{1}{KL(\rho,\rho^{*})},$ (3)
where we note that the confidence measure will reduce as the KL divergence between $\rho$ and $\rho^{*}$ increases. The the range of a confidence measure, which is inversely proportional to entropy, is bounded between $[0,\infty]$. Ideally, we would like to have the confidence measure as high as possible, which requires us to reduce $KL(\rho,\rho^{*})$. We first expand and note that
$KL(\rho,\rho^{*})=H(\rho)-\sum_{j}\rho(y^{j})\log(\rho^{*}(y^{j})),$ (4)
where $H(\rho):=\sum_{j}\rho(y^{j})\log(\rho(y^{j}))$.
Notably, a higher confidence score—which indicates a lower KL divergence between $\rho$ and $\rho^{*}$—is desirable, as it signifies a closer alignment with the ground truth. In this work, we empirically show that adding the vocal cue in the input helps to reduce the $\text{KL}(\rho,\rho^{*})$ by plotting the confidence measure equation 3 for robotic navigation tasks in Section 5.
@@ 5.4 Winning Rate
Besides indicating low variances and multiple choice distribution change, we also calculate the winning rate. We compare the human annotations for each DNIA data clip with the highest log-probability choices generated by the model to ascertain whether the choices and reasoning processes generated by the LLMs are in harmony with human conceptual understanding.
Winning $\text{Winning rate}=\frac{N_{\text{succ}}}{N_{\text{total}}}$ (5)
where $N_{\text{succ}}$ represents the number of instances where the LLM selects an option that aligns with the DNIA dataset’s human annotations, and $N_{\text{total}}$ denotes the total number of audio clips that prompt the LLM. Essentially, the winning rate is analogous to a success rate, ranging from 0% to 100%, with a higher winning rate indicating better performance.
The winning rate reported in our paper is calculated over the entire DNIA dataset, which consists of 500 audio clips: 285 for LU and 215 for VU.
The results, as detailed in Table 3, show the limitation of current LLMs in explaining vocal uncertainty from only textual input and the effectiveness of our audio augmentation approach. Given that the audio clips originated from human speakers, the most reliable measure of their trustworthiness was human perception, hence the incorporation of a user study to gauge the winning rate with which the language model identified uncertainties in the audio clips and selected the appropriate next-step action.
@@ 7 Conclusions and Open Questions
Questions Regarding Setup. While explicit disambiguation, such as asking an additional question, could indeed clarify instructions, it is not always feasible in real-time interactions where immediacy is crucial. Our approach is designed to allow robots to infer potential uncertainties based on vocal cues, enabling them to act more autonomously and adaptively in dynamic environments. This is particularly important in real-world scenarios, such as theme parks or shopping malls, where visitors are often unfamiliar with the environment. As first-time visitors, they might provide uncertain or incomplete answers due to their lack of knowledge. In such settings, a robot cannot rely on repeatedly asking for clarification to find a confident response. Instead, our method allows the robot to navigate uncertainty and make informed decisions without the need for continuous human interaction. While it is true that some existing studies focus on having robots directly re-ask humans to align uncertainty, such as in Ren et al. (2023), our work serves as a complementary approach. We specifically target the robot’s ability to infer uncertainties based on vocal cues alone, without the need for explicit clarification. This focus on vocal cue-based uncertainty detection is crucial in enhancing the robot’s ability to detect genuine uncertainty, leading to more intuitive and seamless human-robot interactions. We acknowledge that disfluencies can arise from external factors, such as distractions, but our method is designed to help robots discern when such disfluencies are indicative of genuine uncertainty, thereby improving their decision-making process without relying on continuous human input.
@@
Broader Impact Statement
This research has the potential to advance our understanding of vocal cue interpretation, particularly in fields such as large language models, audio uncertainty quantification, and human-robot interaction. However, it is crucial to acknowledge and address potential biases that may arise in interpreting vocal cues. These biases could stem from factors such as gender, accent, or cultural background, which may influence how vocal signals are perceived and understood. While our approach is effective, our approach is restricted by the capability of current state-of-the-art Large-Language Models. Therefore, it should be applied thoughtfully and is more appropriate for less sensitive situations to reduce the risk of biases impacting downstream tasks.
Acknowledgments
We are deeply grateful to Bufan Xiao, Xuanting “Fina” Zhou, Yutong “Chloe” Wu, Tongjia Zhang, Muqing Tang, Linqing Liu, Xindi Tang, Phu Pham, and Hrishikesh Viswanath for their invaluable contributions to the DNIA dataset collection and their assistance in revising this paper.
@@ Appendix E DNIA Dataset Details and Samples
For VU clips, we ensure no semantic uncertainty signals, such as uncertain-related words (i.e., ‘maybe,’ ‘I guess’), speech repairs, and hesitations. We believe those are common affective characteristics associated with human speech. However, those affective cues are not prominent in LU clips. As our position states that affective cues (duration, pitch, loudness) are associated with speech uncertainty levels in robot navigation tasks, we believe the setup of our dataset is comprehensive enough to show that vocal cues are good predictors. DNIA dataset is in GitHub repo: https://github.com/beyond-text/DNIA-dataset. The dataset will come with textual annotations (grounded labels) as well as Whisper-generated transcription to save resources for other researchers who can reproduce the result.
Details on human choices for 500 DNIA audio clips are as follows:
Table 7: Human Choices Distribution
A B C D E
Total 66 154 95 96 89
LU 40 90 52 53 50
VU 26 64 43 43 39
The user-study website is shown in Figure 7. The instructions emphasized careful listening and unbiased selection, with a constraint of one minute per clip to simulate real-world decision-making time pressures. We recruited a diverse group of 20 participants, varying in gender and dialect backgrounds, to ensure a broad range of perspectives.
From the dataset’s ground truth scores, it’s evident that the majority of correct choices fall within the ‘B’,‘C’,‘D’,‘E’ options. This aligns with the results in Figure 4, further demonstrating that our method achieves higher accuracy than the baselines.
@@ Appendix F Pitch and Loudness Shifts Detection
1:audio $\mathcal{V}$, $\theta_{l}$ (loudness threshold), $\theta_{p}$ (pitch threshold)
2:Timestamps of maximum loudness change ($t_{l}$) and pitch shift ($t_{p}$)
3:$L\leftarrow\emptyset$, $P\leftarrow\emptyset$triangleright Loudness and Pitch changes
4:$l_{\text{prev}},p_{\text{prev}}\leftarrow$ loudness and pitch at $[t_{0},t_{1}]$
5:for each second $t_{i}$ in $\mathcal{V}$ do
6:     Define segment as $[t_{i},t_{i+1}]$
7:     $l_{t_{i+1}}\leftarrow$ avg_loudness(segment)
8:     $p_{t_{i+1}}\leftarrow$ calculate_pitch(segment)
9:     if $|l_{t_{i+1}}-l_{\text{prev}}|>\theta_{l}$ then
10:         Append ($t_{i+1}$, $|l_{t_{i+1}}-l_{\text{prev}}|$) to $L$
11:     end if
12:     if $|p_{t_{i+1}}-p_{\text{prev}}|>\theta_{p}$ then
13:         Append ($t_{i+1}$, $|p_{t_{i+1}}-p_{\text{prev}}|$) to $P$
14:     end if
15:     $l_{\text{prev}}\leftarrow l_{t_{i+1}}$, $p_{\text{prev}}\leftarrow p_{t_{i+1}}$
16:end for
17:$t_{l}\leftarrow\text{argmax}_{t}\left\{\delta_{l_{t}}\ \middle|\ (t,\delta_{l_{t}})\in L\right\}$
18:$t_{p}\leftarrow\text{argmax}_{t}\left\{\delta_{p_{t}}\ \middle|\ (t,\delta_{p_{t}})\in P\right\}$
Appendix G Sensitivity Study to Loudness and Pitch Threshold
We conducted a threshold experiment on loudness and pitch to demonstrate the effectiveness of our method. With loudness levels set at 0.03 and 0.05, and pitch set at 300 Hz and 500 Hz respectively, the winning rate of our method consistently outperformed the baseline methods by a significant margin. Specifically, compared to the "Transcription Only LLM," which had a winning rate of 22.16% (total), 21.75% (Language Uncertainty, LU), and 22.79% (Vocal Uncertainty, VU), and the "With Chain-of-Thought Reasoning" method, which achieved 49.30% (total), 58.60% (Language Uncertainty, LU), and 36.74% (Vocal Uncertainty, VU), our method showed superior performance (These baseline scores are also shown in Table 3. While we acknowledge that the robustness may not be consistent across all scenarios, the results indicate that the proposed method can be effectively generalized to diverse environments.
Table 8: Sensitivity Study
Condition Total LU VU
Loudness 0.03, Pitch 300 Hz 0.704 0.688 0.726
Loudness 0.05, Pitch 300 Hz 0.614 0.600 0.633
Loudness 0.03, Pitch 500 Hz 0.616 0.593 0.647
Loudness 0.05, Pitch 500 Hz 0.602 0.596 0.609
Appendix H User Interface for Human Evaluation on DNIA Dataset